# Timely Information Updating for Mobile Devices Without and With ML Advice

### v1

@@ 23    $x(t)\leftarrow x(t)+\frac{1}{C_{M}}\sum_{\tau=T_{i}}^{t}x(\tau)+\frac{1}{\theta C_{M}}$;
Combining the lower bound in Lemma 15 with the matching achievability scheme in Alg. 3, we establish that the optimal consistency–robustness trade-off is characterized by the pair of $(\lambda e^{\lambda/R})/(e^{\lambda/R}-1)$ and $(e^{\lambda/R})/(e^{\lambda/R}-1)$. This matches the result for the classic online TCP with ML [bamas2020primal] when $R=1$ (without cost variation).
Note that in many prior studies on ML-augmented online algorithms (e.g., [bamas2020primal, liu2025learning]), the consistency approaches 1 as the trust parameter $\lambda\to 0$, i.e., setting $\lambda\to 0$ forces the algorithm to rely entirely on the ML advice. In contrast, in our setting the consistency approaches $R$ as $\lambda\to 0$, which exceeds 1 whenever $R>1$. This difference is explained as follows. The robustness becomes unbounded as $\lambda\to 0$. Because this trade-off is optimal, the robustness is infinite whenever the consistency falls below $R$. This implies that robustness guarantees collapse even when an online algorithm follows ML advice only partially (so that its consistency remains below $R$). Thus, in our setting, taking $\lambda\to 0$ does not force the algorithm to rely fully on the ML advice. Instead, it identifies the limit in which the algorithm becomes as consistent with the ML advice as possible while still maintaining a finite robustness guarantee.
@@ 29             $x(t)\leftarrow x(t)+\frac{1}{C_{M}}\sum_{\tau=T_{i}}^{t}x(\tau)+\frac{1}{\theta C_{M}}$;
    $\frac{1}{C_{M}}\left(\sum_{\tau=T_{i}}^{t}x(\tau)\bigg|_{\text{after the activation}}\right)+\frac{1}{\theta C_{M}}.$
    However, the paired $z_{i}(t^{\prime})$ was already set to be $\sum_{\tau=T_{i}}^{t^{\prime}}x(\tau)$ in a slot $t^{\prime}\leq t$. Because $x(\tau)$ does not change (for all possible $\tau$) over the virtual OFF period until slot $t$, we have
@@
    $C(t)\left[\frac{1}{C_{M}}\left(\sum_{\tau=T_{i}}^{t}x(\tau)\Big|_{\text{after the activation}}\right)+\frac{1}{\theta C_{M}}\right]$
@@
    $+\left(\sum_{\tau=T_{i}}^{t}x(\tau)\bigg|_{\text{after the activation}}\right)-\left(\sum_{\tau=T_{i}}^{t}x(\tau)\bigg|_{\text{start of slot~$t$}}\right).$
@@
    $\left(\sum_{\tau=T_{i}}^{t}x(\tau)\big|_{\text{after the activation}}\right)-\left(\sum_{\tau=T_{i}}^{t}x(\tau)\bigg|_{\text{start of slot~$t$}}\right)$
@@
    If a virtual packet $i$ arrives before $\hat{t}$ but does not activate in iteration $t^{\prime}$ of Line 4 in slot $t$, then because the paired $z_{i}(t^{\prime})$ was still set to $1-\sum_{\tau=T_{i}}^{t^{\prime}}x(\tau)$ in slot $t^{\prime}\leq t$, it also contributes at most 11 to the objective value in that iteration.
@@ VIII Numerical studies
We adopt a setting similar to [hsu2019scheduling], which derived optimal offline scheduling policies under stationary assumptions. We simulate a horizon of $T=10000$ slots. Update opportunities follow a Bernoulli process with rate $0.7$. To model update costs with memory, we use a two-state Markov chain (as in the Gilbert–Elliott model) with states $\mathsf{L}$ (low cost) and $\mathsf{H}$ (high cost). Let $S(t)\in\{\mathsf{L},\mathsf{H}\}$ denote the state in slot $t$, and let $\mathrm{tr}_{p}$ and $\mathrm{tr}_{q}$ be the transition probabilities from $\mathsf{L}$ to $\mathsf{H}$ and from $\mathsf{H}$ to $\mathsf{L}$, respectively.
Following [hsu2019scheduling], an optimal offline scheduling policy in the stationary setting can be characterized by two age thresholds: if $S(t)=\mathsf{H}$, the device transmits when the age reaches a threshold $T_{\mathsf{H}}$; if $S(t)=\mathsf{L}$, it does when the age reaches a threshold $T_{\mathsf{L}}$. We compute the optimal pair of $T_{\mathsf{H}}$ and $T_{\mathsf{L}}$ via exhaustive search to minimize the total cost in Eq. (2).
@@ VIII-B Nonlinear Aging Function
In this section, we consider a nonlinear aging function similar to that in [kosta2017age]. Specifically, if $x$ slots have elapsed since the most recent update, then the age of information in the current slot is given by $\lfloor e^{0.3x}\rfloor-1$. Throughout this section, we fix $R=2$.
@@ X Acknowledgments
We thank the authors of [liu2025learning] for pointing out mistakes in our earlier preliminary work [tseng2019online]. This research was supported by the National Science and Technology Council, Taiwan, under Grant No. 110-2221-E-305-008-MY3 and 113-2628-E-305-001-MY3.
@@ Appendix B Proof of Theorem 6
Next, let $J(k)$ denote the increment of the objective value in Eq. (5a) by Alg. 1, according to the activations of all virtual packets that arrive in period $k$. This includes the increments of $z_{i}(t)$ in Line 1 and of $x(t)$ in Line 1, for all virtual packet $i$ with $t_{k-1}+1\leq T_{i}\leq t_{k}$ and for all slots $t$. The objective value computed by Alg. 1 is then $\sum_{k=1}^{n+1}J(k)$. We analyze $J(k)$ in two cases.
@@ Appendix D Proof of Lemma 9
This transformation does not alter the optimal solution. We redefine $C(t)/C_{m}$ as the new (normalized) clearing cost. Under the given instance, the transformation yields a clearing cost of 11 in slot 11, and a clearing cost of $R$ in all subsequent slots. Moreover, the term $1/C_{m}$ can be interpreted as the cost of holding a virtual packet for a slot of duration $1/C_{m}$, under the convention that holding a virtual packet for one unit time incurs a unit cost. As $C_{m}\to\infty$, the slot length approaches zero, and the problem transitions to a continuous-time scheduling model, similar to prior studies [bamas2020primal]. In this continuous-time setting, we assume that time starts at 0. The clearing cost becomes $C(0)=1$ and $C(t)=R$ for all $t>0$. Moreover, we consider a time horizon $T\in(0,1]$, which is unknown to the device .
@@
Comparing with the right-hand side cTcT, we conclude that
@@ Appendix F Proof of Lemma 12
for all $i\in P(t)$. If $N_{f}=0$, then Eq. (14) (after the inequality) equals 11. Next, we show that Eq. (14) is nondecreasing in $N_{f}$. Differentiating it with respect to $N_{f}$ gives
@@
where $(a)$ sets $N_{f}=0$ (since $(1+(1/C_{M}))^{(C_{m}-N_{f})/\lambda}$ decreases in $N_{f}$). From [bamas2020primal, Page 27], we can write $(1+(1/C_{M}))^{C_{M}}=e^{x}$ for some $x\in(0,1)$. Let $x^{\prime}=x/R\in(0,1)$. Then, Eq. (15) becomes
@@
which is known to be nonnegative [bamas2020primal, Page 27]. Hence, Eq. (14) is nondecreasing as its derivative is nonnegative, proving the claim.
@@ Appendix G Proof of Theorem 13
for some $x^{\prime}\in(0,1)$. This expression is known to be nonnegative [bamas2020primal, Page 27]. Hence, Eq. (17) is nondecreasing in $N_{f}(k)$, and its maximum is following attained at $N_{f}(k)=C_{m}+1$,
@@ Appendix I Proof of Lemma 15
Setting $T=1$ and assuming perfect ML advice (updating at time 0), the ML advice incurs a cost of 11, while the online algorithm incurs a cost of $\int_{0}^{1}(R+t)\,p(t)\,dt$. By $\big(\lambda e^{\lambda/R}\big)/(e^{\lambda/R}-1)$-consistency, we have
@@
    $\frac{1}{C_{M}}\left(\sum_{\tau=T_{i}}^{t}x(\tau)\bigg|_{\text{after the activation}}\right)+\frac{1}{\theta C_{M}}.$
    However, the paired $z_{i}(t^{\prime})$ was already set to be $\sum_{\tau=T_{i}}^{t^{\prime}}x(\tau)$ in a slot $t^{\prime}\leq t$. Because $x(\tau)$ does not change (for all possible $\tau$) over the virtual OFF period until slot $t$, we have
@@
    $C(t)\left[\frac{1}{C_{M}}\left(\sum_{\tau=T_{i}}^{t}x(\tau)\Big|_{\text{after the activation}}\right)+\frac{1}{\theta C_{M}}\right]$
@@
    $+\left(\sum_{\tau=T_{i}}^{t}x(\tau)\bigg|_{\text{after the activation}}\right)-\left(\sum_{\tau=T_{i}}^{t}x(\tau)\bigg|_{\text{start of slot~$t$}}\right).$ (25)
@@
    $\left(\sum_{\tau=T_{i}}^{t}x(\tau)\big|_{\text{after the activation}}\right)-\left(\sum_{\tau=T_{i}}^{t}x(\tau)\bigg|_{\text{start of slot~$t$}}\right)$
@@
    If a virtual packet $i$ arrives before $\hat{t}$ but does not activate in iteration $t^{\prime}$ of Line 4, then because the paired $z_{i}(t^{\prime})$ was still set to $1-\sum_{\tau=T_{i}}^{t^{\prime}}x(\tau)$ in slot $t^{\prime}\leq t$, it contributes at most 11 to the objective value in that iteration.
@@
    If $p(1)=1$: Set $T=1$, $\Delta A(1)=0$, and $U(1)=1$. In this case, the online algorithm incurs cost $C_{m}$, while the offline optimum chooses not to update and incurs total cost 11. Hence, the competitive ratio is $C_{m}$, which again diverges as $C_{m}\to\infty$.
@@
    If $p(1)=1$: Set $T=1$ and $U(1)=1$. In this case, the online algorithm incurs cost $C_{m}$, while the offline optimum chooses not to update and incurs total cost 11. Hence, the competitive ratio is $C_{m}$, which again diverges as $C_{m}\to\infty$.

### v2

@@ 23    $x(t)\leftarrow x(t)+\frac{1}{C_{M}}\sum_{\tau=T_{i}}^{t}x(\tau)+\frac{1}{\theta C_{M}}$;
Combining the lower bound in Lemma 15 with the matching achievability scheme in Alg. 3, we establish that the optimal consistency–robustness trade-off is characterized by the pair of $(\lambda e^{\lambda/R})/(e^{\lambda/R}-1)$ and $(e^{\lambda/R})/(e^{\lambda/R}-1)$. This matches the result for the classic online TCP with ML [3] when $R=1$ (without cost variation).
Note that in many prior studies on ML-augmented online algorithms (e.g., [3, 16]), the consistency approaches 1 as the trust parameter $\lambda\to 0$, i.e., setting $\lambda\to 0$ forces the algorithm to rely entirely on the ML advice. In contrast, in our setting the consistency approaches $R$ as $\lambda\to 0$, which exceeds 1 whenever $R>1$. This difference is explained as follows. The robustness becomes unbounded as $\lambda\to 0$. Because this trade-off is optimal, the robustness is infinite whenever the consistency falls below $R$. This implies that robustness guarantees collapse even when an online algorithm follows ML advice only partially (so that its consistency remains below $R$). Thus, in our setting, taking $\lambda\to 0$ does not force the algorithm to rely fully on the ML advice. Instead, it identifies the limit in which the algorithm becomes as consistent with the ML advice as possible while still maintaining a finite robustness guarantee.
@@ 29             $x(t)\leftarrow x(t)+\frac{1}{C_{M}}\sum_{\tau=T_{i}}^{t}x(\tau)+\frac{1}{\theta C_{M}}$;
    $\frac{1}{C_{M}}\left(\sum_{\tau=T_{i}}^{t}x(\tau)\bigg|_{\text{before the activation}}\right)+\frac{1}{\theta C_{M}}.$
    However, the paired $z_{i}(t^{\prime})$ was already set to be $1-\sum_{\tau=T_{i}}^{t^{\prime}}x(\tau)$ in a slot $t^{\prime}\leq t$. Because $x(\tau)$ does not change (for all possible $\tau$) over the virtual OFF period until slot $t$, we have
@@
    $C(t)\left[\frac{1}{C_{M}}\left(\sum_{\tau=T_{i}}^{t}x(\tau)\Big|_{\text{before the activation}}\right)+\frac{1}{\theta C_{M}}\right]$
@@
    $+\left(\sum_{\tau=T_{i}}^{t}x(\tau)\bigg|_{\text{before the activation}}\right)-\left(\sum_{\tau=T_{i}}^{t}x(\tau)\bigg|_{\text{start of slot~$t$}}\right).$
@@
    $\left(\sum_{\tau=T_{i}}^{t}x(\tau)\big|_{\text{before the activation}}\right)-\left(\sum_{\tau=T_{i}}^{t}x(\tau)\bigg|_{\text{start of slot~$t$}}\right)$
@@
    If a virtual packet $i$ arrives before $\hat{t}$, meets the condition in Line 4 in slot $t^{\prime}\leq t$, but does not activate in iteration $t^{\prime}$ of Line 4 in slot $t$, then because the paired $z_{i}(t^{\prime})$ was still set to $1-\sum_{\tau=T_{i}}^{t^{\prime}}x(\tau)$ in slot $t^{\prime}$, it also contributes at most 11 to the objective value in that iteration.
@@ VIII Numerical studies
We adopt a setting similar to [8], which derived optimal offline scheduling policies under stationary assumptions. We simulate a horizon of $T=10000$ slots. Update opportunities follow a Bernoulli process with rate $0.7$. To model update costs with memory, we use a two-state Markov chain (as in the Gilbert–Elliott model) with states $\mathsf{L}$ (low cost) and $\mathsf{H}$ (high cost). Let $S(t)\in\{\mathsf{L},\mathsf{H}\}$ denote the state in slot $t$, and let $\mathrm{tr}_{p}$ and $\mathrm{tr}_{q}$ be the transition probabilities from $\mathsf{L}$ to $\mathsf{H}$ and from $\mathsf{H}$ to $\mathsf{L}$, respectively.
Following [8], an optimal offline scheduling policy in the stationary setting can be characterized by two age thresholds: if $S(t)=\mathsf{H}$, the device transmits when the age reaches a threshold $T_{\mathsf{H}}$; if $S(t)=\mathsf{L}$, it does when the age reaches a threshold $T_{\mathsf{L}}$. We compute the optimal pair of $T_{\mathsf{H}}$ and $T_{\mathsf{L}}$ via exhaustive search to minimize the total cost in Eq. (2).
@@ VIII-B Nonlinear aging function
In this section, we consider a nonlinear aging function similar to that in [12]. Specifically, if $x$ slots have elapsed since the most recent update, then the age of information in the current slot is given by $\lfloor e^{0.3x}\rfloor-1$. Throughout this section, we fix $R=2$.
@@ X Acknowledgments
We thank the authors of [16] for pointing out mistakes in our earlier preliminary work [26]. This research was supported by the National Science and Technology Council, Taiwan, under Grant No. 110-2221-E-305-008-MY3 and 113-2628-E-305-001-MY3.
@@ Appendix B Proof of Theorem 6
Next, let $J(k)$ denote the increment of the objective value in Eq. (5a) by Alg. 1, according to the activations of all virtual packets that arrive in period $k$. This includes the increments of $z_{i}(t)$ in Line 1 and of $x(t)$ in Line 1, for all virtual packets $i$ with $t_{k-1}+1\leq T_{i}\leq t_{k}$ and for all slots $t$. The objective value computed by Alg. 1 is then $\sum_{k=1}^{n+1}J(k)$. We analyze $J(k)$ in two cases.
@@ Appendix D Proof of Lemma 9
This transformation does not alter the optimal solution. We redefine $C(t)/C_{m}$ as the new (normalized) clearing cost. Under the given instance, the transformation yields a clearing cost of $1$ in slot $1$, and a clearing cost of $R$ in all subsequent slots. Moreover, the term $1/C_{m}$ can be interpreted as the cost of holding a virtual packet for a slot of duration $1/C_{m}$, under the convention that holding a virtual packet for one unit time incurs a unit cost. As $C_{m}\to\infty$, the slot length approaches zero, and the problem transitions to a continuous-time scheduling model, similar to prior studies [3]. In this continuous-time setting, we assume that time starts at 0. The clearing cost becomes $C(0)=1$ and $C(t)=R$ for all $t>0$. Moreover, we consider a time horizon $T\in(0,1]$, which is unknown to the device .
@@
Comparing with the right-hand side $cT$, we conclude that
@@ Appendix F Proof of Lemma 12
for all $i\in P(t)$. If $N_{f}=0$, then Eq. (14) (after the inequality) equals $1$. Next, we show that Eq. (14) is nondecreasing in $N_{f}$. Differentiating it with respect to $N_{f}$ gives
@@
where $(a)$ sets $N_{f}=0$ (since $(1+(1/C_{M}))^{(C_{m}-N_{f})/\lambda}$ decreases in $N_{f}$). From [3, Page 27], we can write $(1+(1/C_{M}))^{C_{M}}=e^{x}$ for some $x\in(0,1)$. Let $x^{\prime}=x/R\in(0,1)$. Then, Eq. (15) becomes
@@
which is known to be nonnegative [3, Page 27]. Hence, Eq. (14) is nondecreasing as its derivative is nonnegative, proving the claim.
@@ Appendix G Proof of Theorem 13
for some $x^{\prime}\in(0,1)$. This expression is known to be nonnegative [3, Page 27]. Hence, Eq. (17) is nondecreasing in $N_{f}(k)$, and its maximum is following attained at $N_{f}(k)=C_{m}+1$,
@@ Appendix I Proof of Lemma 15
Setting $T=1$ and assuming perfect ML advice (updating at time 0), the ML advice incurs a cost of $1$, while the online algorithm incurs a cost of $\int_{0}^{1}(R+t)\,p(t)\,dt$. By $\big(\lambda e^{\lambda/R}\big)/(e^{\lambda/R}-1)$-consistency, we have
@@
    $\frac{1}{C_{M}}\left(\sum_{\tau=T_{i}}^{t}x(\tau)\bigg|_{\text{before the activation}}\right)+\frac{1}{\theta C_{M}}.$
    However, the paired $z_{i}(t^{\prime})$ was already set to be $1-\sum_{\tau=T_{i}}^{t^{\prime}}x(\tau)$ in a slot $t^{\prime}\leq t$. Because $x(\tau)$ does not change (for all possible $\tau$) over the virtual OFF period until slot $t$, we have
@@
    $C(t)\left[\frac{1}{C_{M}}\left(\sum_{\tau=T_{i}}^{t}x(\tau)\Big|_{\text{before the activation}}\right)+\frac{1}{\theta C_{M}}\right]$
@@
    $+\left(\sum_{\tau=T_{i}}^{t}x(\tau)\bigg|_{\text{before the activation}}\right)-\left(\sum_{\tau=T_{i}}^{t}x(\tau)\bigg|_{\text{start of slot~$t$}}\right).$ (25)
@@
    $\left(\sum_{\tau=T_{i}}^{t}x(\tau)\big|_{\text{before the activation}}\right)-\left(\sum_{\tau=T_{i}}^{t}x(\tau)\bigg|_{\text{start of slot~$t$}}\right)$
@@
    If a virtual packet $i$ arrives before $\hat{t}$, meets the condition in Line 4 in slot $t^{\prime}\leq t$, but does not activate in iteration $t^{\prime}$ of Line 4 in slot $t$, then because the paired $z_{i}(t^{\prime})$ was still set to $1-\sum_{\tau=T_{i}}^{t^{\prime}}x(\tau)$ in slot $t^{\prime}$, it contributes at most 11 to the objective value in that iteration.
@@
    If $p(1)=1$: Set $T=1$, $\Delta A(1)=0$, and $U(1)=1$. In this case, the online algorithm incurs cost $C_{m}$, while the offline optimum chooses not to update and incurs total cost $1$. Hence, the competitive ratio is $C_{m}$, which again diverges as $C_{m}\to\infty$.
@@
    If $p(1)=1$: Set $T=1$ and $U(1)=1$. In this case, the online algorithm incurs cost $C_{m}$, while the offline optimum chooses not to update and incurs total cost $1$. Hence, the competitive ratio is $C_{m}$, which again diverges as $C_{m}\to\infty$.